AshtynThatDrummerGirl:

Question 1 A 1.0 L sample of gas at 1.5 atm and 273K has the volume and temperature raised to 5.0 L and 298 K, What pressure is the sample under? 0.33 atm 3.1 atm 0.27 atm 3.6 atm Question 2 Gases will behave like an ideal gas at____? STP absolute zero room temperature any temperature and pressure Question 3  Calculate the moles of a 5.0 L sample of gas at 101.3kPa and 298 K. 0.20 mol 20.7 mol 1.0 mol 1.20 mol

AshtynThatDrummerGirl:

@BenLindquist @lilyrn @NAM @Nnesha @pandasurvive @PirateKing1999 @quickstudent @ThisGirlPretty @Ultrilliam

Nnesha:

\[\huge\rm \frac{ P_1V_1 }{ T_1 }=\frac{P_2V_2}{T_2}\] \[\rm P_1 =1.5 ~atm ~~,~~~~~V_2=1.0~L~~~,~~~~~~T_1=273 K\]\[\rm P_2 =?(find) ~~,~~~~~V_2=5.0~L~~~,~~~~~~T_1=298 K\] plugin the values, solve for for P_2

Nnesha:

use ideal gas law for the 3rd one \[\large\rm PV=nRT\] R is constant \[\large\rm R=8.31 \frac{KPa \cdot L}{mole\cdot K}\]

AshtynThatDrummerGirl:

Thank you so much!!!

Nnesha:

yw